1. Assume the 2 out 5 shipments coming from a particular supplier do not arrive on time. What is the probability that among 8 shipments which this supplier sends,
a. Exactly 3 will not arrive on time?
b. Less than 3 will not arrive on time?
c. Exactly 6 will arrive on time?
2. The probability that a certain production process will produce a defective part is 0.20. Find the probability that a lot of 12 parts will contain exactly 5 defective parts.
a. Exactly 5 defectives?
b. 3 or 4 defective parts?
3. The average number of points which a PBA players makes per quarter of a game is 6.5, find the probability that in the first quarter of the next game, he will make 7 to 10 points?
THe IQs of 600 applicants of a certain college are approximately normally distributed with a mean of 115 and a standard deviation of 12. If the college requires an IQ of at least 95, how many of thse students will be rejected on this basis regardless of their other qualifications?

Answer by stanbon(48568)

1. Assume the 2 out 5 shipments coming from a particular supplier do not arrive on time. What is the probability that among 8 shipments which this supplier sends,
1. Assume the 2 out 5 shipments coming from a particular supplier do not arrive on time. What is the probability that among 8 shipments which this supplier sends,
Prob(arrive on time)= 3/5; Prob(do not arrive on time) = 2/5
a. Exactly 3 will not arrive on time?
=8C3 (2/5)^3 (3/5)^5 = 0.27869...
b. Less than 3 will not arrive on time?
=Prob(0 will not)+Prob(1 will not)+Prob(2 will not)
=0.3153
c. Exactly 6 will arrive on time?
0.20901888...
2. The probability that a certain production process will produce a defective part is 0.20. Find the probability that a lot of 12 parts will contain exactly 5 defective parts.
Prob(defective part)=0.2 ; Prob(not defective)=0.8
a. Exactly 5 defectives?
0.0531...
b. 3 or 4 defective parts?
0.36909875...
3. The average number of points which a PBA players makes per quarter of a game is 6.5, find the probability that in the first quarter of the next game, he will make 7 to 10 points?
Convert the 7 and the 10 to z-scores; then find the probability that lies
between those z values. To do this you have the mean but you also need
the standard deviation of his scoring.
--------------------------------------------------
THe IQs of 600 applicants of a certain college are approximately normally distributed with a mean of 115 and a standard deviation of 12. If the college requires an IQ of at least 95, how many of thse students will be rejected on this basis regardless of their other qualifications?
Do a z-score on 95 using mean 115 and standard deviation of 12. Find the probability that lies below that z-score. Ans: 0.048
Cheers,
Stan H.